# publications.bib

### v3

@@ 4.3. Verification of additional assumptions
Throughout this section, we assume $(\mathbf{G},X)$ is a PEL datum of type $A$ or $C$ such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of simple groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions, and that Assumption 1.10 holds.
@@
(Theorem 1.14) By Corollary 3.17, the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a $G(\mathbb{Q}_{p})\times W_{E}$-equivariant filtration with graded pieces isomorphic to $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$. The cohomology of Igusa varieties $V_{b}$ and the global Shimura variety is admissible [Zha, Proposition 8.21], so we can apply the results of the previous section to them. We let $\mathfrak{m}$ be a generic maximal ideal of the spherical Hecke algebra, and consider the localization
$(j_{1}^{*}T_{\mu}j_{b!}(V_{b}))_{\phi_{\mathfrak{m}}}[-d](-\frac{d}{2}).$
This defines a filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$. The filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ considered above comes from applying $(-)_{\phi_{\mathfrak{m}}}$ to $R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))$ viewed as a $G(\mathbb{Q}_{p})$-representation. Using Corollary 4.25, we see that these graded pieces are also isomorphic to
$(j_{1}^{*}T_{\mu}j_{b*}(V_{b}))_{\phi_{\mathfrak{m}}}[-d](-\frac{d}{2})$
via the natural transformation $j_{b!}\rightarrow j_{b*}$ and are trivial for $b\in B(G,\mu)_{\mathrm{un}}$. However, using Lemma 3.14, this implies that the natural map
$R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))_{\phi_{\mathfrak{m}}}\rightarrow R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}],i_{b*}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))_{\phi_{\mathfrak{m}}}$
is an isomorphism, using Remark 3.15. Therefore, we see that the edge maps in the excision spectral sequence actually degenerate after applying $(-)_{\phi_{\mathfrak{m}}}$, giving us a direct sum decomposition
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\simeq\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})_{\phi_{\mathfrak{m}}}.$
By applying $R\Gamma(K_{p}^{\mathrm{hs}},-)$ and invoking Lemma 4.2 (3), we obtain that
$R\Gamma(K_{p}^{\mathrm{hs}},R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p}},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}})\simeq R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p}K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$
has a filtration with graded pieces isomorphic to
$R\Gamma(K_{p}^{\mathrm{hs}},j_{1}^{*}T_{\mu}j_{b!}(V_{b}))_{\mathfrak{m}}[-d](-\frac{d}{2}).$
Just as in the proof of Corollary 3.17, we can rewrite this as
$(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}V_{b})[2d_{b}],$
@@
(Theorem 1.16) We recall, by Proposition 3.7, that $V_{b}$ is a complex of smooth $J_{b}(\mathbb{Q}_{p})$-representations concentrated in degree $\leq d_{b}$. It follows that we have an inclusion $\bigoplus_{b\in B(G,\mu)}j_{b!}(V_{b})_{\phi_{\mathfrak{m}}}\in\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0,\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$, using Proposition A.5. Corollary 4.24 implies that
@@
is concentrated in degrees $0\leq i\leq d$. By applying Poincaré duality at finite level and Corollary A.7, this allows us to conclude that the non-compactly supported cohomology
@@
localized at $\phi_{\mathfrak{m}}^{\vee}$ is concentrated in degrees $d\leq i\leq 2d$, where we define this to be the colimit over the non-compactly supported cohomology of finite levels (cf. Remark 3.1). Moreover, we note that generic is preserved under the Chevalley involution, since it just exchanges the role of positive and negative roots. It therefore follows that
@@
is also concentrated in degrees $\geq d$, but this isomorphic to
@@
by Lemma 4.2 (3). This establishes Theorem 1.16, by applying Poincaré duality to the Shimura variety at finite level again. ∎
@@
We would like to obtain the main theorem for Shimura varities of non-PEL type, especially Hilbert-Siegel modular varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$ or $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$). We will show this in a more general setup, as follows.
Let $(\mathbf{G},X)$, $(\mathbf{G}_{2},X_{2})$ be a pair of abelian type Shimura data such that $\mathbf{G},\mathbf{G}_{2}$ are centrally isogenous, and we have an isomorphism of derived subgroups
$\mathbf{G}^{\mathrm{der}}\xrightarrow{\sim}\mathbf{G}_{2}^{\mathrm{der}},$
as well as adjoint quotients. Consider the associated Shimura varieties $\mathrm{Sh}(\mathbf{G},X)_{K}$ and $\mathrm{Sh}(\mathbf{G}_{2},X_{2})_{K_{2}}$, where we choose the level $K,K_{2}$ such that the level at $p$, satisfies that we have an equality $K_{p}\cap G^{\mathrm{der}}(\mathbb{Q}_{p})=K_{2,p}\cap G_{2}^{\mathrm{der}}(\mathbb{Q}_{p})$.
We now assume $K_{p}$ and $K_{2,p}$ are both hyperspecial. Observe that this implies that $K^{\prime}_{p}=K_{p}\cap G^{\mathrm{der}}(\mathbb{Q}_{p})$ is also hyperspecial. By the Satake isomorphism, we have an isomorphism of $\overline{\mathbb{F}}_{\ell}$-algebras
@@
Fix a connected component $X^{+}\subset X$. This also fixes a $X^{+}_{2}\subset X_{2}$, and an isomorphism $X^{+}\simeq X^{+}_{2}$, since $\mathbf{G},\mathbf{G}_{2}$ have isomorphic adjoint quotients. For any compact open subgroup $K\subset\mathbf{G}(\mathbb{A}_{f})$, we let $\mathrm{Sh}^{+}(\mathbf{G},X)_{K}$ be the geometrically connected component which is the image of $X^{+}\times 1$. Moreover, we will let
Note that since all the transition morphisms are finite étale and hence affine, $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}$ is also a qcqs scheme by [stacks-project, Lemma 01YX].
Since $\mathbf{G},\mathbf{G}_{2}$ have isomorphic derived subgroups, this implies that we have an isomorphism of geometric connected components
$\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}\simeq\mathrm{Sh}^{+}(\mathbf{G}_{2},X^{\prime})_{K_{2,p}}.$
Moreover, we see that the action of $H^{\prime}_{K^{\prime}_{p}}$ on $\mathrm{Sh}(\mathbf{G},X)_{K_{p}}$ preserves the geometric connected component, since we see that $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}$ is simply the connected Shimura variety associated to $(\mathbf{G}^{\mathrm{der}},X^{+})$. Indeed, we see that the set of $\mathbb{C}$-points of $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}$ is given by
$\mathbf{G}^{\mathrm{der}}(\mathbb{Q})^{(p),\mathrm{cl}}_{+}\backslash X^{+}\times(\mathbf{G}^{\mathrm{der}}(\mathbb{A}_{f}^{p})),$
where $\mathbf{G}^{\mathrm{der}}(\mathbb{Q})^{(p),\mathrm{cl}}_{+}$ denotes the closure in $\mathbf{G}^{\mathrm{der}}(\mathbb{A}_{f}^{p})$ of $\mathbf{G}^{\mathrm{der}}(\mathbb{R})_{+}\cap\mathbf{G}^{\mathrm{der}}(\mathbb{Q})\cap K_{p}$, and $\mathbf{G}^{\mathrm{der}}(\mathbb{R})_{+}$ denotes the preimage of the neutral connected component $\mathbf{G}^{\mathrm{ad}}(\mathbb{R})^{+}$ in $\mathbf{G}^{\mathrm{der}}(\mathbb{R})$ under the quotient map.
Following the description of the connected components of Shimura varities from [Del, §2], and using the notation of [Kis, §3.3] we see that there exist groups $\mathscr{A}(\mathbf{G})^{\circ},\mathscr{A}(\mathbf{G}_{2}),\mathscr{A}(\mathbf{G})$ such that we have $\mathbf{G}(\mathbb{A}_{f})$ (resp. $\mathbf{G}_{2}(\mathbb{A}_{f})$) equivariant isomorphisms
$\mathrm{Sh}(\mathbf{G},X)_{K_{p}}\simeq[\mathscr{A}(\mathbf{G})\times\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}/\mathscr{A}(\mathbf{G})^{\circ}]$
and
$\mathrm{Sh}(\mathbf{G}_{2},X_{2})_{K_{2,p}}\simeq[\mathscr{A}(\mathbf{G}_{2})\times\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}/\mathscr{A}(\mathbf{G})^{\circ}].$
Observe that, since $\mathscr{A}(\mathbf{G})^{\circ}=\mathscr{A}(\mathbf{G}^{\mathrm{der}})^{\circ}$ is a subgroup of $\mathscr{A}(\mathbf{G}_{2})$, the Shimura variety $\mathrm{Sh}(\mathbf{G}_{2},X_{2})_{K_{2,p}}$ is simply an (infinite) union of copies of $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}$. Moreover, we see that the action of $H^{\prime}_{K^{\prime}_{p}}$ on the right-hand side of the above isomorphism is given by the action on $\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}}$. In particular, we observe that
$H^{i}(\mathrm{Sh}^{+}(\mathbf{G},X)_{K_{p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$
vanishes if and only if
$H^{i}(\mathrm{Sh}(\mathbf{G}_{2},X)_{K_{2,p}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m^{\prime}}}$
does as well. We thus have the following proposition.
Proposition 5.1.
Suppose that $(\mathbf{G},X)$ is of PEL type A or C satisfying the conditions in Theorem 1.16. Then, Conjecture 1.2 also holds for $(\mathbf{G}_{2},X_{2})$.
Proof.
Since Conjecture 1.2 is true for $(\mathbf{G},X)$, we will first show that a maximal ideal $\mathfrak{m}$ of $H_{K_{p}}$ is generic if and only if the maximal ideal $\mathfrak{m}^{\prime}$ of $H^{\prime}_{K^{\prime}_{p}}$ is generic. To see this, we will reformulate this in terms of $L$-parameters. This is equivalent to showing that an $L$-parameter
$\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$
is generic if and only if the composition $\phi^{\prime}$ with the map $g:\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\rightarrow\phantom{}^{L}T^{\prime}(\overline{\mathbb{F}}_{\ell})$ induced by the inclusion of tori $T^{\prime}\hookrightarrow T$ is generic. (Here, $T^{\prime}=G^{\mathrm{der}}\cap T$). This follows from the observation that any coroot $\alpha$ factors through $G^{\mathrm{der}}$, and hence the composition $\alpha\circ\phi$ is equal to $\alpha\circ\phi^{\prime}$.
Now, we consider the limit
$\mathrm{Sh}(\mathbf{G},X):=\mathop{\rm lim}\nolimits\limits_{{\begin{subarray}{c}\leftarrow\\ K^{p}\end{subarray}}}\mathrm{Sh}(\mathbf{G},X)_{K^{p}}.$
Note that since all schemes appearing in the limit are qcqs, by [stacks-project, Theorem 09YQ] we have an isomorphism of cohomology groups
$H^{i}(\mathrm{Sh}(\mathbf{G},X),\overline{\mathbb{F}}_{\ell})\simeq\mathop{\rm lim}\nolimits\limits_{{\begin{subarray}{c}\rightarrow\\ K^{p}\end{subarray}}}H^{i}(\mathrm{Sh}(\mathbf{G},X)_{K^{p}},\overline{\mathbb{F}}_{\ell}).$
We now have the following lemma.
@@
Let $G^{\prime}\rightarrow G$ be a map inducing an isomorphism on adjoint groups with $g:\phantom{}^{L}G\rightarrow\phantom{}^{L}G^{\prime}$, the induced map on dual groups. For $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ a L-parameter and $A$ an admissible complex of $G(\mathbb{Q}_{p})$-modules, there is a natural isomorphism of $G^{\prime}(\mathbb{Q}_{p})$-modules
$(A|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}\simeq\bigoplus_{\begin{subarray}{c}\phi\\ \phi^{\prime}=g\circ\phi\end{subarray}}A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})},$
with notation as in Corollary 4.3.

### v5

@@ 5. The Proof of the Main Theorems
We can now reap the fruit of our work in the previous sections and finally prove our main results. We start with Theorem 1.17.
Theorem 5.1.
Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions and let $\phi_{T}$ be a generic toral parameter with associated semi-simple L-parameter $\phi$ of $G$. Then the complex $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,C},\overline{\mathbb{F}}_{\ell})_{\phi}$ breaks up as a direct sum
$\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\phi}\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma_{c-\partial}(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell}))[2d_{b}]$
of $G(\mathbb{Q}_{p})$-modules.
@@
By Corollary 3.22, the complex $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})$ has a $G(\mathbb{Q}_{p})$-equivariant filtration with graded pieces isomorphic to $j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})$. The cohomology of Igusa varieties $V_{b}$ and the global Shimura variety is admissible as a complex (For the $V_{b}$, we can write $V_{b}$ as the fiber of the natural excision map $R\Gamma(\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})\rightarrow R\Gamma(\partial\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})$ as in the proof of Corollary 3.5 and then analogously to [68, Proposition 8.21] show that $R\Gamma(\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\partial\mathrm{Ig}^{b,*},\overline{\mathbb{F}}_{\ell})$ are admissible, since the condition of admissibility is preserved under finite limits, the claim follows), and therefore it follows that $j_{b!}(V_{b})\in\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, and so we can apply the results of the previous section to them. We consider the localization
$(j_{1}^{*}T_{\mu}j_{b!}(V_{b}))_{\phi}[-d](-\frac{d}{2}).$
This defines a filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi}$. The filtration on $R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ considered above comes from applying $(-)_{\phi}$ to $R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))$ viewed as a $G(\mathbb{Q}_{p})$-representation. Using Corollary 4.29, we see that these graded pieces are also isomorphic to
$(j_{1}^{*}T_{\mu}Rj_{b*}(V_{b}))_{\phi}[-d](-\frac{d}{2})$
via the natural transformation $j_{b!}\rightarrow Rj_{b*}$ and are trivial for $b\notin B(G,\mu)_{\mathrm{un}}$. However, using Lemma 3.18, this implies that the natural map
$R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b!}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))_{\phi_{\mathfrak{m}}}\rightarrow R\Gamma([\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}],i_{b*}i_{b}^{*}(R\pi_{\mathrm{HT!}}(\overline{\mathbb{F}}_{\ell})))_{\phi_{\mathfrak{m}}}$
is an isomorphism (See Remark 3.19). Therefore, we see that the edge maps in the excision spectral sequence actually degenerate after applying $(-)_{\phi}$, giving us a direct sum decomposition
$R\Gamma_{c}(\mathcal{S}(\mathbf{G},X)_{K^{p},C},\overline{\mathbb{F}}_{\ell})_{\phi}\simeq\bigoplus_{b\in B(G,\mu)_{\mathrm{un}}}j_{1}^{*}T_{\mu}j_{b!}(V_{b})[-d](-\frac{d}{2})_{\phi},$
@@
We now turn our attention to Theorem 1.8.
Theorem 5.2.
Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions, and let $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ be a generic maximal ideal with associated semi-simple L-parameter $\phi_{\mathfrak{m}}$. Then, for a level $K=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology of $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
@@
We recall, by Proposition 3.6, that $V_{b}$ is a complex of smooth $J_{b}(\mathbb{Q}_{p})$-representations concentrated in degree $\leq d_{b}$. It follows that we have $\bigoplus_{b\in B(G,\mu)}j_{b!}(V_{b})_{\phi_{\mathfrak{m}}}\in\phantom{}^{\mathrm{p}}\mathrm{D}^{\leq 0,\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$, using Proposition A.5. Corollary 4.27 implies that
@@
is concentrated in degrees $0\leq i\leq d$. By applying Poincaré duality at the pro-$p$ finite levels and Corollary A.7, this allows us to conclude that the non-compactly supported cohomology
@@
localized at $\phi_{\mathfrak{m}}^{\vee}$ is concentrated in degrees $d\leq i\leq 2d$, where we define this to be the colimit over the non-compactly supported cohomology of finite levels (cf. Remark 3.7). Moreover, we note that genericity is preserved under taking duals and the same is true for regularity. It therefore follows that
@@
is also concentrated in degrees $\geq d$, but this is isomorphic to
@@
by Lemma 4.3 (3). This establishes Theorem 1.8, by applying Poincaré duality to the Shimura variety at finite level $K_{p}^{\mathrm{hs}}$. ∎
We now turn our attention to the abelian type case.
@@
We would like to obtain the main Theorem for some Shimura varities of non-PEL type, especially Hilbert-Siegel modular varieties (attached to $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GL}_{2}$ or $\mathrm{Res}_{F/\mathbb{Q}}\mathrm{GSp}_{4}$). We will show this in a more general setup, as follows.
Let $(\mathbf{G},X)$, $(\mathbf{G}_{1},X_{1})$ be a pair of abelian type Shimura data such that $\mathbf{G},\mathbf{G}_{1}$ are centrally isogenous, and we have an isomorphism of derived subgroups
$\mathbf{G}^{\mathrm{der}}\xrightarrow{\sim}\mathbf{G}_{1}^{\mathrm{der}},$
(and hence of adjoint quotients). We further assume that both groups $G,G_{1}$ are unramified (observe that we can always arrange this, up to changing the prime $p$). Consider the associated Shimura varieties $\mathrm{Sh}(\mathbf{G},X)_{K}$ and $\mathrm{Sh}(\mathbf{G}_{1},X_{1})_{K_{1}}$, where we will choose the level $K,K_{1}$ such that the level at $p$ satisfies that we have an equality $K_{p}\cap G^{\mathrm{der}}(\mathbb{Q}_{p})=K_{1p}\cap G_{1}^{\mathrm{der}}(\mathbb{Q}_{p})$. In fact, we can and will take $K_{p}$ and $K_{1p}$ hyperspecial and hence $K^{\prime}_{p}=K_{p}\cap G^{\mathrm{der}}(\mathbb{Q}_{p})$ is also hyperspecial. This is possible because we observe that the group $G^{\mathrm{der}}$ is hence also unramified, and we have a natural map of Bruhat-Tits buildings
$\mathcal{B}(G^{\mathrm{der}},\breve{\mathbb{Q}}_{p})\rightarrow\mathcal{B}(G,\breve{\mathbb{Q}}_{p})$
which sends a hyperspecial point on $\mathcal{B}(G^{\mathrm{der}},\breve{\mathbb{Q}}_{p})$ to a hyperspecial point, so by fixing a choice of hyperspecial point of $\mathcal{B}(G^{\mathrm{der}},\breve{\mathbb{Q}}_{p})$ we can obtain hyperspecial subgroups $K_{p}^{\prime},K_{p},K_{1p}$ as desired.
Now, observe that by the Satake isomorphism, we have an isomorphism of $\overline{\mathbb{F}}_{\ell}$-algebras
@@
We first want to understand how this passage to the derived subgroup works for localization at various $L$-parameters $\phi$. We have the following Lemma:
@@
Let $G^{\prime}\rightarrow G$ be a map inducing an isomorphism on adjoint groups with $g:\phantom{}^{L}G\rightarrow\phantom{}^{L}G^{\prime}$, the induced map on dual groups. For $\phi:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ a L-parameter and $A$ an admissible complex of $G(\mathbb{Q}_{p})$-modules, then there is a natural isomorphism of $G^{\prime}(\mathbb{Q}_{p})$-modules
$(A|_{G^{\prime}(\mathbb{Q}_{p})})_{\phi^{\prime}}\simeq\bigoplus_{\begin{subarray}{c}\phi\\ \phi^{\prime}=g\circ\phi\end{subarray}}A_{\phi}|_{G^{\prime}(\mathbb{Q}_{p})}.$
where the LHS is the composite $\mathrm{D}(G^{\prime}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})\xrightarrow{j_{1!}}\mathrm{D}(\mathrm{Bun}_{G^{\prime}},\overline{\mathbb{F}}_{\ell})\xrightarrow{(-)_{\phi^{\prime}}}\mathrm{D}(\mathrm{Bun}_{G^{\prime}},\overline{\mathbb{F}}_{\ell})_{\phi^{\prime}}$ applied to $A|_{G^{\prime}(\mathbb{Q}_{p})}$.